OpenStudy (naveenbhatia1312):

Assuming all variables are positive, use properties of logarithms to write the expression as a sum or difference of logarithms or multiples of logarithms.

OpenStudy (naveenbhatia1312):

log\[\log_{2}7x \]

OpenStudy (mathstudent55):

Rule of logarithms: \(\Large \log ab = \log a + \log b\)

OpenStudy (naveenbhatia1312):

would it be log7 + log x?

OpenStudy (naveenbhatia1312):

@mathstudent55

OpenStudy (naveenbhatia1312):

where woulf the ^2 go?

OpenStudy (mathstudent55):

correct

OpenStudy (naveenbhatia1312):

would it be log 1 + log 1 or log 2 + log? @mathstudent55

OpenStudy (naveenbhatia1312):

@Zale101

OpenStudy (mathstudent55):

The 2 is the base of the log, so every log on the right side must have that base. \(\log_2 7x = \log_2 7 + \log_2 x\)